how many gallons of a 90% antifreeze solution must be mixed with 90 gallons of 15% antifreeze to get a mixture that is 80% antifreeze?
Answer by mananth(16946)

------ Percent ---------------- quantity
Antifreeze 15 ----------------90
Antifreeze 90 ----------------x
Total 80 ---------------- 90 + x

15*90+90*x=80(90+x)
1350+90x=7200+80x
90x-80x =7200-1350
10x=5850
/10
x=585 gallons 90 % Antifreeze